Hydra Report - May 2002: System work & Eng Results
Detector, dewar and Arcon work, Ramon Galvez H., June 2002
- Leak tested.
- Determine gasification rate = 1.8 SCFH
- Determine As Is Thermal Radiation Load = 3.8 Watts
- New vacuum valve installed.
CCD & Arcon
- Dark Current = 1.43 e/pix/hr ; Conditions = Bin 2x2, Gain 1, all sorrounding rooms and dome lights ON.
- Dark Current = 0.88 e/pix/hr ; Conditions = Bin 2x2, Gain 1, all lights OFF.
- New ADC bd prepared + tested ( s/n4 new batch). Got no more missing codes around 15k ADU.
- Video original and spare bd. Noise comparison : both at 3e.
- Not a single crash. Neither during exposure tests nor while downloading.
- A new LN2 refill and solidification device was designed. I had the good advices from Brooke on this. Tested in the lab and, along with Daniel, on the telescope.
- Cryo Status : 24hrs hold time, one daily refill. Cryo system ‘tuned’; as a result the refill is done without necessity of vacuum valve control.
Updated on June 15, 2021, 7:49 am